# RS-RAG: Bridging Remote Sensing Imagery and Comprehensive Knowledge with a Multi-Modal Dataset and Retrieval-Augmented Generation Model

### v1

@@ III-A Data Processing
TABLE I: World Knowledge Fields and Descriptions for Landmark in the RSWK dataset.
@@
TABLE II: Remote Sensing Domain Knowledge Fields, Descriptions, and Source for Landmark in the RSWK dataset.
Field Name Description Source
Albedo Surface albedo, representing the reflectivity of the Earth’s surface MERRA-2 (M2T1NXRAD)
Albnirdf Near-infrared diffuse surface albedo MERRA-2 (M2T1NXRAD)
Albnirdr Near-infrared direct surface albedo MERRA-2 (M2T1NXRAD)
Albvisdf Visible light diffuse albedo MERRA-2 (M2T1NXRAD)
Albvisdr Visible light direct albedo MERRA-2 (M2T1NXRAD)
Emis Surface emissivity, related to thermal radiation MERRA-2 (M2T1NXRAD)
Evland Evaporation land ($\mathrm{kg}/\mathrm{m}^{2}/\mathrm{s}$) MERRA-2 (M2T1NXLND)
Gwetprof Profile soil moisture averaged over depth MERRA-2 (M2T1NXLND)
LC_type1 Land cover type (IGBP classification) MODIS (MCD12Q1.061)
LC_type2 Land cover type (UMD classification) MODIS (MCD12Q1.061)
LC_type3 Land cover type (LAI classification) MODIS (MCD12Q1.061)
LC_type4 Land cover type (BGC classification) MODIS (MCD12Q1.061)
LC_type5 Land cover type (Plant Functional Types classification) MODIS (MCD12Q1.061)
LST_Day_1km Daytime land surface temperature (°C) MODIS (MOD11A1.061)
LST_Night_1km Nighttime land surface temperature (°C) MODIS (MOD11A1.061)
Prectotland Total land precipitation ($\mathrm{kg}/\mathrm{m}^{2}/\mathrm{s}$) MERRA-2 (M2T1NXLND)
Saa Sun Azimuth Angle (°) HLSL30 (HLS-2)
Slp Sea level pressure (Pa) MERRA-2 (M2T1NXSLV)
Smland Snowmelt flux land ($\mathrm{kg}/\mathrm{m}^{2}/\mathrm{s}$) MERRA-2 (M2T1NXLND)
Speed Surface wind speed ($\mathrm{m}/\mathrm{s}$) MERRA-2 (M2T1NXFLX)
Sr_b1 Band 1 (ultra blue, coastal aerosol) surface reflectance Landsat 9
Sr_b2 Band 2 (blue) surface reflectance Landsat 9
Sr_b3 Band 3 (green) surface reflectance Landsat 9
Sr_b4 Band 4 (red) surface reflectance Landsat 9
Sr_b5 Band 5 (near infrared) surface reflectance Landsat 9
Sr_b6 Band 6 (shortwave infrared 1) surface reflectance Landsat 9
Sr_b7 Band 7 (shortwave infrared 2) surface reflectance Landsat 9
Ts Surface skin temperature (°C) MERRA-2 (M2T1NXSLV)
Ulml Surface eastward wind ($\mathrm{m}/\mathrm{s}$) MERRA-2 (M2T1NXFLX)
Vaa View Azimuth Angle (°) HLSL30 (HLS-2)
Vlml Surface northward wind speed ($\mathrm{m}/\mathrm{s}$) MERRA-2 (M2T1NXFLX)
Vza View Zenith Angle (°) HLSL30 (HLS-2)
Mean_2m_air_temperature Average air temperature at 2m height (°C) ERA5 (Daily Aggregates)
Mean_sea_level_pressure Mean sea level pressure (Pa) ERA5 Daily Aggregates
Surface_pressure Surface pressure (Pa) ERA5 (Daily Aggregates)
Total_precipitation Total precipitation (m) ERA5 (Daily Aggregates)
@@ III-B Data Statistics
Through the above data construction pipeline, the RSWK dataset successfully collected a total of 14,141 landmark instances from 175 countries, each accompanied by high-resolution satellite imagery, domain-specific remote sensing attributes, and structured world knowledge descriptions. The global spatial distribution of the landmarks is visualized in Fig. 2(a), demonstrating wide coverage across all major continents. Notably, the dataset includes landmarks from a diverse range of regions, effectively covering most major countries worldwide. To provide a more detailed view of the dataset composition, we present the distribution of landmarks across the top 100 countries sorted by landmark count in Fig. 2(b, left). It can be observed that countries such as the United States, the United Kingdom, China, and Japan contribute the largest number of landmarks. This reflects both the cultural richness and the degree of documentation available for landmarks in these regions. In addition, Fig. 2(b, right) shows the frequency distribution of the top 15 landmark categories, including parks, museums, natural scenery, universities, and historic sites. The category distribution is relatively balanced, highlighting the diverse types of landmarks captured in the dataset and ensuring a rich set of semantic concepts for downstream tasks. To provide an intuitive example of how the data is structured, we present the Sydney Opera House in Fig. 2(c) as a representative landmark. The figure illustrates the three core components of each RSWK entry: the satellite imagery, the domain knowledge (e.g., albedo, emissivity, land cover type, and meteorological variables), and the structured world knowledge (e.g., historical background, architectural characteristics, and cultural significance). This tri-modal representation demonstrates the depth and richness of the dataset, supporting a wide range of geospatial understanding and vision-language reasoning tasks.
@@ IV-A Problem Formulation
To address this limitation, we extend the standard VLM formulation by incorporating external, query-relevant knowledge. Specifically, we adopt a Retrieval-Augmented Generation (RAG) framework, wherein a set of top-$k$ relevant knowledge snippets $\mathbf{R}$ is retrieved from an external corpus based on the multimodal similarity between the input image–text pair $(q_{I},q_{T})$ and the knowledge index. The retrieved context $\mathbf{R}$ is then integrated with the original inputs to guide the response generation process. Formally, the objective becomes:
@@ IV-C Knowledge Retrieval and Response Generation
Figure 4: Qualitative results of image captioning on remote sensing imagery of the Great Seto Bridge. Text in red indicates the recognized landmark name; purple highlights retrieved world knowledge, such as historical, geographic, or cultural facts; and green denotes domain-specific knowledge, including spectral indices, land cover, and ALBEDO values.
TABLE III: Performance comparison between baseline models and our proposed RS-RAG variants on the image captioning task using a subset of the RSWK dataset.
@@
InternVL2.5-8B [29] 0.427 0.289 0.218 0.166 0.222 0.275 0.013
Janus-Pro-7B [30] 0.370 0.245 0.183 0.139 0.197 0.246 0.007
Qwen2.5-VL-7B [31] 0.349 0.223 0.161 0.121 0.185 0.229 0.004
Llama3.2-Vision-11B [32] 0.448 0.303 0.228 0.173 0.225 0.286 0.014
RS-RAG-7B 0.409 0.263 0.193 0.144 0.206 0.240 0.065
RS-RAG-11B 0.470 0.360 0.307 0.266 0.276 0.322 0.027
TABLE IV: Performance comparison between baseline models and our proposed RS-RAG variants on the image classfication task using a subset of the RSWK dataset.
Model Overall Airport Amusement Beach Bridge Casino Church Gov. Bldg. Historic Mansion Museum Park Stadium Theater Tower University
@@
InternVL2.5-8B [29] 0.329 0.959 0.250 0.667 0.385 0.000 0.000 0.152 0.158 0.023 0.000 0.672 0.790 0.039 0.053 0.654
Janus-Pro-7B [30] 0.269 0.959 0.500 0.778 0.308 0.000 0.480 0.485 0.061 0.000 0.000 0.373 0.632 0.000 0.053 0.161
Qwen2.5-VL-7B [31] 0.309 0.959 0.300 0.556 0.615 0.091 0.112 0.364 0.094 0.023 0.016 0.709 0.790 0.039 0.158 0.296
Llama3.2-Vision-11B [32] 0.340 0.959 0.300 0.667 0.462 0.091 0.020 0.061 0.052 0.047 0.278 0.746 0.816 0.039 0.053 0.568
RS-RAG-7B 0.659 0.959 0.800 0.667 0.692 0.636 0.551 0.576 0.474 0.419 0.349 0.866 0.868 0.346 0.368 0.765
RS-RAG-11B 0.842 0.980 0.900 0.944 0.769 0.727 0.820 0.849 0.655 0.512 0.714 0.843 0.842 0.769 0.526 0.803
@@ V-A Experimental Setup
All input images are resized to 512 by 512 pixels before being fed into the models. We evaluate several state-of-the-art vision-language models as baselines, including InternVL2.5-Instruct-8B [29], Janus-Pro-7B [30], Qwen-2.5-VL-7B-Instruct [31], and LLaMA-3.2-Vision-11B-Instruct [32]. Building upon Qwen-2.5-VL and LLaMA-3.2-Vision, we further develop our retrieval-augmented variants, named RS-RAG-7B and RS-RAG-11B, respectively. To better adapt these models to the remote sensing domain, we fine-tune them via Low-Rank Adaptation (LoRA) [33] using instruction-following data consisting of 1380 curated image-text pairs. Fine-tuning is conducted for 3 epochs with a batch size of 1, using the Adamw optimizer and an initial learning rate of $1\times 10^{-4}$. To systematically evaluate instruction-tuned vision-language models in the remote sensing domain, we curate a new benchmark consisting of task-specific subsets for image captioning (348 samples), image classification (910 samples), and visual question answering (300 samples). All experiments are performed using 3 NVIDIA RTX A6000 GPUs, each with 48 GB of memory.
To comprehensively evaluate model performance, we adopt a set of standard metrics tailored to each task. For image captioning and visual question answering, we report BLEU 1, BLEU 2, BLEU 3, BLEU 4, METEOR, ROUGE L, and CIDEr to measure the fluency, relevance, and informativeness of generated text. For the classification task, we evaluate both overall accuracy and per class accuracy to reflect model performance across diverse scene categories.
@@ V-B Results on Image Captioning Task
Table III summarizes the performance of baseline and proposed models on the image captioning task over remote sensing imagery, evaluated using a comprehensive set of metrics including BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr. Among the baseline models, LLaMA3.2-Vision-11B achieves the strongest performance, attaining a BLEU-4 score of 0.173 and METEOR of 0.225, highlighting its relative strength in generating syntactically and semantically coherent captions. However, it still exhibits limitations in domain-specific understanding and factual richness due to its lack of grounding in external knowledge. In contrast, both of our proposed retrieval-augmented models, RS-RAG-7B and RS-RAG-11B, demonstrate consistent improvements across all evaluation metrics. Notably, RS-RAG-11B outperforms all baselines with a BLEU-4 of 0.266 and METEOR of 0.276, representing absolute gains of +0.093 and +0.051 over the strongest baseline. It also achieves the highest ROUGE-L score of 0.322, indicating improved phrase-level overlap with human-written descriptions. While RS-RAG-7B uses a smaller backbone, it still yields a significant performance boost, achieving the best CIDEr score of 0.065, which suggests enhanced alignment with human judgment of caption quality.
To better understand the effectiveness of our RS-RAG framework, we conduct a qualitative comparison of image captioning outputs generated by baseline models and our RS-RAG model, as shown in Fig. 4. The selected case depicts the Great Seto Bridge, a large-scale landmark in Japan. Baseline models, such as InternVL2.5-8B and Qwen2.5-VL-7B, either misidentify the landmark (e.g., predicting the Rainbow Bridge or Akashi Kaikyo Bridge) or generate overly generic descriptions without domain-grounded facts. Although LLaMA3.2-Vision-11B produces a correct name, it lacks detailed evidence from the image or environmental context. Notably, our proposed RS-RAG-7B and RS-RAG-11B models generate accurate, entity-aware, and knowledge-rich descriptions that correctly identify the Great Seto Bridge and provide factual details about its coordinates, construction history, and structural type. Furthermore, the descriptions include domain-specific knowledge (e.g., surface albedo, emissivity, and land cover classification), demonstrating the ability of our models to integrate retrieved remote sensing metadata. The use of world knowledge, such as historical background and engineering facts, further enhances the semantic richness and correctness of the generated captions. These results highlight the benefits of retrieval-augmented generation in grounding VLM outputs with both domain and general-purpose knowledge.

### v2

@@ III-A Data Collection and Processing
TABLE I: Remote Sensing Domain Knowledge Fields, Descriptions, Temporal Resolution, and Source for Landmark in the RSWK dataset.
Field Name Description Temporal Resolution Source
Albedo Surface albedo, representing the reflectivity of the Earth’s surface Annual, Monthly MERRA-2 (M2T1NXRAD)
Albnirdf Near-infrared diffuse surface albedo Annual, Monthly MERRA-2 (M2T1NXRAD)
Albnirdr Near-infrared direct surface albedo Annual, Monthly MERRA-2 (M2T1NXRAD)
Albvisdf Visible light diffuse albedo Annual, Monthly MERRA-2 (M2T1NXRAD)
Albvisdr Visible light direct albedo Annual, Monthly MERRA-2 (M2T1NXRAD)
Emis Surface emissivity, related to thermal radiation Annual, Monthly MERRA-2 (M2T1NXRAD)
Evland Evaporation land (mm) Annual, Monthly MERRA-2 (M2T1NXLND)
Gwetprof Profile soil moisture averaged over depth Annual, Monthly MERRA-2 (M2T1NXLND)
LC_type1 Land cover type (IGBP classification) Annual MODIS (MCD12Q1.061)
LC_type2 Land cover type (UMD classification) Annual MODIS (MCD12Q1.061)
LC_type3 Land cover type (LAI classification) Annual MODIS (MCD12Q1.061)
LC_type4 Land cover type (BGC classification) Annual MODIS (MCD12Q1.061)
LC_type5 Land cover type (Plant Functional Types classification) Annual MODIS (MCD12Q1.061)
LST_Day_1km Daytime land surface temperature (°C) Annual, Monthly MODIS (MOD11A1.061)
LST_Night_1km Nighttime land surface temperature (°C) Annual, Monthly MODIS (MOD11A1.061)
Prectotland Total land precipitation (mm) Annual, Monthly MERRA-2 (M2T1NXLND)
Saa Sun Azimuth Angle (°) Annual HLSL30 (HLS-2)
Slp Sea level pressure (Pa) Annual, Monthly MERRA-2 (M2T1NXSLV)
Smland Snowmelt flux land (mm) Annual, Monthly MERRA-2 (M2T1NXLND)
Speed Surface wind speed ($\mathrm{m}/\mathrm{s}$) Annual, Monthly MERRA-2 (M2T1NXFLX)
Sr_b1 Band 1 (ultra blue, coastal aerosol) surface reflectance Annual, Monthly Landsat 9
Sr_b2 Band 2 (blue) surface reflectance Annual, Monthly Landsat 9
Sr_b3 Band 3 (green) surface reflectance Annual, Monthly Landsat 9
Sr_b4 Band 4 (red) surface reflectance Annual, Monthly Landsat 9
Sr_b5 Band 5 (near infrared) surface reflectance Annual, Monthly Landsat 9
Sr_b6 Band 6 (shortwave infrared 1) surface reflectance Annual, Monthly Landsat 9
Sr_b7 Band 7 (shortwave infrared 2) surface reflectance Annual, Monthly Landsat 9
Ts Surface skin temperature (°C) Annual, Monthly MERRA-2 (M2T1NXSLV)
Ulml Surface eastward wind ($\mathrm{m}/\mathrm{s}$) Annual, Monthly MERRA-2 (M2T1NXFLX)
Vaa View Azimuth Angle (°) Annual HLSL30 (HLS-2)
Vlml Surface northward wind speed ($\mathrm{m}/\mathrm{s}$) Annual, Monthly MERRA-2 (M2T1NXFLX)
Vza View Zenith Angle (°) Annual HLSL30 (HLS-2)
Mean_2m_air_temperature Average air temperature at 2m height (°C) Annual, Monthly ERA5 (Daily Aggregates)
Mean_sea_level_pressure Mean sea level pressure (Pa) Annual, Monthly ERA5 Daily Aggregates
Surface_pressure Surface pressure (Pa) Annual, Monthly ERA5 (Daily Aggregates)
Total_precipitation Total precipitation (mm) Annual, Monthly ERA5 (Daily Aggregates)
NDVI Normalized Difference Vegetation Index Annual, Monthly Landsat 9
NDWI Normalized Difference Water Index Annual, Monthly Landsat 9
NDBI Normalized Difference Building Index Annual, Monthly Landsat 9
TABLE II: World Knowledge Fields and Descriptions for Landmark in the RSWK dataset.
@@ III-B Data Statistics
Through the above data construction pipeline, the RSWK dataset successfully collected a total of 14,820 landmark instances from 184 countries, each accompanied by high-resolution satellite imagery, structured remote sensing knowledge, and comprehensive world knowledge descriptions. The imagery is categorized into 16 representative scene classes, including Airport, Amusement Park, Beach, Bridge, Casino, Church, Government Building, Historic Site, Mansion, Museum, Park, Stadium, Theater, Tower, University, and Forest. This categorization captures both urban and natural scenes, enabling the dataset to support a wide range of geospatial vision-language tasks. The remote sensing knowledge module comprises 39 expert-derived variables, as detailed in Table I, including surface reflectance, land surface temperature, albedo, precipitation, and soil moisture. In parallel, each landmark is enriched with 14 structured world knowledge fields, as shown in Table II, capturing descriptive and contextual information such as Historical Background, Cultural Significance, Architectural Characteristics, and Primary Function.
@@
The global spatial distribution of the landmarks is visualized in Fig.2(a), demonstrating wide coverage across all major continents. Notably, the dataset includes landmarks from a diverse range of regions, effectively covering most major countries worldwide. To provide a more detailed view of the dataset composition, we present the distribution of landmarks across the top 50 countries sorted by landmark count in Fig.2(b, left). It can be observed that countries such as the United States, the United Kingdom, China, and Japan contribute the largest number of landmarks. This reflects both the cultural richness and the degree of documentation available for landmarks in these regions. In addition, Fig.2(b, right) shows the frequency distribution of the top 16 landmark categories. The category distribution is relatively balanced, highlighting the diverse types of landmarks captured in the dataset and ensuring a rich set of semantic concepts for downstream tasks. To provide an intuitive example of how the data is structured, we present the Sydney Opera House in Fig.2(c) as a representative landmark. The figure illustrates the three core components of each RSWK entry: the satellite imagery, the domain knowledge, and the structured world knowledge. This tri-modal representation demonstrates the depth and richness of the dataset, supporting a wide range of geospatial understanding and vision-language reasoning tasks.
TABLE III: Overview of dataset statistics across task categories in both full and mini benchmark settings.
Task Category Total (Full) Train (Full) Test (Full) Total (Mini) Train (Mini) Test (Mini)
Image Captioning 14820 11827 2993 7408 5912 1496
Image Classification 14820 11827 2993 4800 3200 1600
VQA VQA_C 22604 18103 4501 11301 9051 2250
VQA_RSK 14820 11827 2993 7408 5912 1496
VQA_WK 14820 11827 2993 7409 5913 1496
III-C Benchmark Setting
To facilitate systematic evaluation of VLMs in the context of remote sensing and world knowledge understanding, we establish a comprehensive benchmark suite based on the RSWK dataset. This benchmark encompasses three core tasks: image captioning, image classification, and VQA. Specifically, the dataset includes 14,820 image captioning samples, each paired with a descriptive sentence that summarizes the visual content of the satellite imagery along with relevant remote sensing expert knowledge and contextual world knowledge. It also includes 14,820 image classification question-answer pairs, where each question queries the semantic category of the depicted scene, and the answer corresponds to a predefined class label. For the VQA task, we construct a total of 52,244 question-answer pairs aimed at evaluating the model’s multimodal reasoning abilities. These VQA instances are further divided into three types: (1) VQA_C: 22,604 comprehensive questions that require holistic understanding of both visual content and combined remote sensing and world knowledge; (2) VQA_RSW: 14,820 remote sensing knowledge questions that involve interpreting scientific and geophysical attributes; and (3) VQA_WK: 14,820 world knowledge questions that focus on encyclopedic or factual information. To support model development and reproducible evaluation, we further split the dataset into 11,827 training samples and 2,993 testing samples. The exact sample counts for each task type are summarized in Table III. Additionally, considering the computational constraints faced by some researchers, we construct a mini-benchmark version comprising approximately 50% of the full data for each task category, with detailed statistics also reported in Table III.
@@ IV-A Problem Formulation
To address this limitation, we extend the standard VLM formulation by incorporating external, query-relevant knowledge. Specifically, we adopt a RAG framework, wherein a set of top-$k$ relevant knowledge snippets $\mathbf{R}$ is retrieved from an external corpus based on the multimodal similarity between the input image–text pair $(q_{I},q_{T})$ and the knowledge index. The retrieved context $\mathbf{R}$ is then integrated with the original inputs to guide the response generation process. Formally, the objective becomes:
@@ IV-C Knowledge Retrieval and Response Generation
Figure 4: Qualitative results of image captioning on remote sensing imagery of the Massachusetts Institute of Technology (MIT). Text in red indicates the identified landmark name, blue highlights visual content directly observable from the image, purple denotes retrieved world knowledge, such as historical, geographic, or cultural facts; and green denotes remote sensing knowledge, including spectral indices, land cover, and ALBEDO values.
TABLE IV: Performance comparison between baseline models and our proposed RS-RAG variants on the image captioning task using the RSWK-Mini dataset.
@@
InternVL2.5 [6] 0.368 0.221 0.154 0.113 0.312 0.325 0.014
Janus-Pro [5] 0.370 0.221 0.154 0.112 0.312 0.324 0.013
Qwen2.5-VL [31] 0.385 0.237 0.169 0.125 0.331 0.344 0.018
Llama3.2-Vision [1] 0.379 0.232 0.164 0.122 0.321 0.338 0.016
RSGPT [10] 0.192 0.097 0.061 0.041 0.144 0.154 0.002
Geochat [13] 0.347 0.201 0.139 0.101 0.291 0.297 0.006
RS-RAG 0.490 0.366 0.300 0.252 0.447 0.476 0.145
TABLE V: Performance comparison between baseline models and our proposed RS-RAG variants on the image classification task using the RSWK-Mini dataset.
Model Overall Airport Amusement Beach Bridge Casino Church Gov. Bldg. Historic Mansion Museum Park Stadium Theater Tower University Forest
@@
InternVL2.5 [6] 0.57 0.91 0.59 0.73 0.76 0.82 0.35 0.42 0.27 0.60 0.16 0.21 0.90 0.45 0.41 0.72 0.86
Janus-Pro [5] 0.58 0.95 0.67 0.62 0.84 0.76 0.40 0.37 0.17 0.64 0.30 0.20 0.90 0.46 0.49 0.64 0.87
Qwen2.5-VL [31] 0.61 0.90 0.66 0.69 0.78 0.77 0.51 0.39 0.23 0.70 0.29 0.32 0.88 0.52 0.59 0.69 0.86
Llama3.2-Vision [1] 0.59 0.93 0.62 0.66 0.77 0.72 0.43 0.39 0.13 0.75 0.41 0.28 0.90 0.48 0.57 0.55 0.90
Rsgpt [10] 0.46 0.91 0.70 0.79 0.75 0.37 0.22 0.18 0.31 0.50 0.06 0.49 0.93 0.06 0.11 0.66 0.39
Geochat [14] 0.60 0.94 0.55 0.75 0.79 0.79 0.54 0.34 0.11 0.65 0.27 0.39 0.88 0.6 0.57 0.64 0.84
RS-RAG 0.79 0.98 0.76 0.77 0.92 0.91 0.81 0.98 0.75 0.92 0.72 0.80 0.99 0.23 0.52 0.74 0.89
@@ V-A Experimental Setup
All input images are resized to 512 by 512 pixels before being fed into the models. We evaluate several state-of-the-art vision-language models as baselines, including InternVL2.5-Instruct-8B [6], Janus-Pro-7B [5], Qwen-2.5-VL-7B-Instruct [31], LLaMA-3.2-Vision-11B-Instruct [1], RSGPT [10], and Geochat [14]. Our proposed RS-RAG model is built upon Qwen-2.5-VL-7B. To better adapt these models to the remote sensing domain, we perform instruction tuning on the RSWK dataset using Low-Rank Adaptation (LoRA) [9]. Fine-tuning is conducted for 3 epochs with a batch size of 1, using the Adamw optimizer and an initial learning rate of $1\times 10^{-4}$. All experiments are performed using 3 NVIDIA RTX A6000 GPUs, each with 48 GB of memory.
To comprehensively evaluate model performance, we adopt a set of standard metrics tailored to each task. For image captioning and visual question answering, we report BLEU-1, BLEU-2, BLEU-3, BLEU-4, METEOR, ROUGE-L, and CIDEr to measure the fluency, relevance, and informativeness of generated text. For the classification task, we evaluate both overall accuracy and per-class accuracy to reflect model performance across diverse scene categories.
@@ V-B Results on Image Captioning Task
After instruction tuning all VLMs, we first evaluate their performance on the image captioning task. Table IV presents a comparative analysis of baseline VLMs and our proposed RS-RAG model on the RSWK-Mini dataset. Across all seven evaluation metrics, BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr, RS-RAG consistently outperforms all baselines by a substantial margin. While Qwen2.5-VL surpasses other baseline models, RS-RAG achieves a BLEU-1 score of 0.490 and a CIDEr score of 0.145, representing improvements of 10.5 and 12.7 points respectively over Qwen2.5-VL. These results highlight the benefit of incorporating structured external knowledge into VLMs, which significantly enhances their capacity to generate accurate, contextually rich, and semantically meaningful captions in the remote sensing domain. The superior performance of RS-RAG underscores the effectiveness of our retrieval-augmented design for complex, knowledge-intensive multimodal reasoning tasks.
To better understand the effectiveness of our RS-RAG framework, we conduct a qualitative comparison of image captioning outputs generated by baseline models and our RS-RAG model, as shown in Fig.4. Specifically, Fig.4 presents a remote sensing image corresponding to the Massachusetts Institute of Technology (MIT). Among all evaluated models, RS-RAG is the only one that correctly identifies the landmark and accurately integrates both factual and scientific domain knowledge into the caption. The output from RS-RAG includes detailed metadata such as the official address, spatial coverage, historical construction timeline, and multiple spectral and physical indicators. This demonstrates RS-RAG’s ability to effectively synthesize geographic information with remote sensing expertise. In contrast, the baseline models exhibit limitations in both semantic accuracy and factual grounding. For instance, InternVL2.5 misidentifies the site as Boston University, GeoChat assigns the location to Cathedral Church, and RSGPT attributes it to the University of Caldas in Colombia. Some baseline models do retrieve domain-relevant attributes, yet the outputs often lack factual coherence due to incorrect or inconsistent landmark recognition. These findings highlight the advantages of RS-RAG in connecting visual content with structured multimodal knowledge representations. The model demonstrates strong capability in supporting knowledge-intensive remote sensing tasks, particularly those requiring both precise semantic understanding and accurate factual alignment.